OpenStudy (sswann222):

What is the relationship between the four fundamental subspaces of a matrix?

OpenStudy (anonymous):

http://www.math.sunysb.edu/~badger/notes/4fund.pdf

OpenStudy (slaaibak):

dimension of column space of A = rank[A] dimension of row space of A = rank[A] dimension of nullspace of A = n - r dimension of nullspace of A^T = m-r

OpenStudy (sswann222):

alright now we have to make sure we find the final connection how they Col(A) = Row(A^t)